OpenStudy (anonymous):

Find the surface area and volume of a sphere having a radius of 4". (Leave answer as an improper fraction.) SA = ? pi sq. in. V = ? pi cu. in.

OpenStudy (mertsj):

A=4pir^2 A=4pi(16)=64pi square inches V=4/3pir^3 V=4/3pi(4)^3=4/3(64)pi = 256/3pi